What is the equation of the line which includes points (1,5) and (-2,-7)?
Answer by lwsshak3(3694)

What is the equation of the line which includes points (1,5) and (-2,-7)?
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
slope,m,=∆y/∆x=change in y/change in x=rise/run
using two given points
m=(-7-5)/(-2-1)=-12/-3=4
equation: y=4x+b
solving for b using one of given points(1,5)
5=4*1+b
b=1
equation of line:
y=4x=1